# Probing Hernquist dark matter with black hole shadows: A comprehensive study of various accretions

### v3

@@ Abstract
The shadow of a black hole is critically dependent on the surrounding accreting matter. We perform a systematic comparative analysis of the observational signatures of a Schwarzschild black hole embedded in a Hernquist dark matter (DM) halo under three distinct accretion scenarios: a geometrically thin disk, a static spherical flow, and an infalling spherical flow. For the thin disk model, we find that direct emission dominates the total observed intensity, while the size and brightness of the lensing and photon rings serve as sensitive probes of the Hernquist DM parameters. From a geometric perspective, the Hernquist DM halo significantly enlarges the photon sphere, resulting in an observable shadow radius approximately $2\%$ to $30\%$ larger than in the vacuum case. Regarding the radiative signatures, the measured intensity profiles, which rely on the particular accretion models, show a general brightness suppression, which is especially affected by the Doppler de-boosting in the infalling scenario. Our results suggest that the size and brightness profile of a black hole shadow provide a valuable theoretical framework for constraining the distribution of dark matter in galactic centers.
@@ I Introduction
To investigate the observational characteristics of black hole images and search for potential indications of novel physics, it is essential to employ simplified accretion models. Schwarzschild black holes were studied for the spherically symmetric accretion model, another accretion scenario, and it was shown that under such accretion settings the shadow has robust properties, where its shape and size are determined by spacetime geometry rather than accretion details Gralla et al. (2019); Tsukamoto et al. (2014); Tsukamoto (2018). Conversely, investigations of optically thin and geometrically thin accretion models show that the appearance of the shadow is strongly related to the position of the emission source Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021b); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Direct emission, lensing rings, and photon rings are the three main components that these models usually appear as in the seen image. These distinct signatures offer a potential means to distinguish GR black holes from other compact objects or black holes in modified gravity theories via observational data Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021b); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Consequently, EHT observations of black hole shadows are crucial for constraining parameters in modified gravity theories and testing fundamental physics Shaikh et al. (2019); Davoudiasl and Denton (2019); Roy and Yajnik (2020).
@@
In this paper, we present a systematic comparative study of the shadow of Schwarzschild-Hernquist black holes with different accretion models. In order to clarify the associated impacts of Hernquist DM halo parameters on the photon sphere and observed brightness, we thoroughly investigated three thin disk accretion models and two spherical accretion models. Our results demonstrate that different accretion regimes produce different shadow images under the same background geometry. Specifically, pictures made up of the shadow, lensing ring, and photon ring are created via thin disk accretion. The shadow and photon ring are spherically symmetric for spherical accretion. A darker shadow than under static conditions results from the extra Doppler effect introduced by falling accretion Li and He (2021a, b); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). We aim to close a gap in the references on Hernquist metrics and accretion models by revealing scaling rules for shadow size and brightness as functions of the Hernquist DM halo distribution. This offers a potential diagnostic framework for constraining Hernquist DM characteristics.
@@ II Light deflection and photon orbit
Here, the Newton constant $G$ has been set to $1$ for simplicity. Naturally, this black hole metric (10) reduces to the Schwarzschild black hole in the limit $\rho_{c}\to 0$ or $r_{s}\to 0$. In Fig.1, we depict the behavior of the metric function $f(r)$ to show how the Hernquist DM halo affects spacetime. The largest root of the equation $f(r_{h})=0$ defines the event horizon radius $r_{h}$. This critical point in the Schwarzschild scenario is located at $r_{h}=2M$. However, the DM halo causes the curve to cross the zero axis at a wider radius, as seen in Fig.1. In particular, a monotonic expansion of the event horizon is induced by an increase in either the core radius $r_{s}$ or the central density $\rho_{c}$. This geometric change indicates that the input from the DM halo may successfully boost the black hole’s.
@@
Figure 3: The behaviour of light trajectories under various $\rho_{c}$ and $r_{s}$ in the polar coordinate system $(r,\varphi)$. The solid black disk and the green dashed circle indicates the event horizon and the photon shpere, respectively. The yellow lines correspond to scattered rays $(b>b_{p})$, the blue lines to captured light $(b<b_{p})$ and the red one means the critical unstable orbit $(b=b_{p})$.
@@ III Observational features of thin disk emission
In astrophysical settings, black holes are typically surrounded by large amounts of accreting matter. This makes the study of shadows cast by Schwarzschild-Hernquist black holes particularly relevant from an observational perspective. In this section we will investigate the shadows and rings of a Schwarzschild-Hernquist black hole surrounded by a disk-like accretion disk that is optically and geometrically thin.
@@ III.1 Direct emission, lensing ring and photon ring
Figure 5: Photon behaviour as a function of impact parameter $b$ in the Schwarzschild-Hernquist black hole. The blue lines represent trajectories contributing to direct emission ($n<3/4$), yellow lines constitute the lensing ring ($3/4<n<5/4$), and red lines mean the photon ring ($n>5/4$). The black disk and the green dashed circle indicates the event horizon and the photon shpere $r_{p}$, respectively.
@@
In favour of the number of orbits $n(b)$, the emission can be categorized into three distinct zones Gralla et al. (2019); Gralla and Lupsasca (2020); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021a, b). These are lensing rings, photon rings, and direct emission rings corresponding to $n<3/4$, $3/4<n<5/4$, and $n>5/4$, respectively. These zones correspond to light rays that intersect the equatorial plane once, twice, or more than twice, respectively Gralla et al. (2019); Gralla and Lupsasca (2020). For various values of the parameters $\rho_{c}$ and $r_{s}$, Fig.4 shows the functional dependence of $n(b)$ on the impact parameter $b$. The domains of direct emissions, lensing rings, and photon rings are shown by different colors. We have indicated the ring and the direct emission locations in Fig.5. The function $n(b)$ diverges as bapproaches the critical value $b_{p}$, indicating the existence of unstable photon orbits, as shown in Fig.4. The curves for larger DM parameters $\rho_{c}$ and $r_{s}$ are shifted to the right, consistent with the increase in $b_{p}$ reported in Tab.1. Tab.2 lists the ranges of the impact parameter bcorresponding to direct emission, the lensing ring, and the photon ring for various values of $\rho_{c}$ and $r_{s}$. It is evident from the data that the range of the impact parameter $b$ for all emission types increases as the parameters increase. Furthermore, the total number of photon orbits $n(b)$ will peak inside a small area when the impact parameter $b$ gets closer to the critical condition $b\sim b_{p}$. The extremely narrow ranges for the photon ring listed in Tab.2 correspond to the thin red band visible in Fig.5. The angular size of the shadow also increases with larger parameters. However, this enlargement does not directly translate to an increase in the brightness of the ring; the observed intensity depends on the complex interplay between the emission profile and the light bending.
@@ III.3 Observational features
To compute the observational appearance of the thin disk, we employ three phenomenological emissivity models Wang et al. (2022); Yang et al. (2023); Li and He (2021b); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). It is important to note that these models are simplified representations and are strategically chosen to represent a diverse range of plausible astrophysical scenarios from efficient, geometrically thin disks to inefficient, thick flows Meng et al. (2023); Gan et al. (2021); Guo et al. (2022); Chen et al. (2022). This approach allows us to systematically explore how the observational signatures of the Hernquist DM halo depend on the spatial distribution of the emitting matter. While the precise quantitative features of the image, such as peak brightness, are model-dependent, we expect the qualitative findings, particularly the scaling of the shadow and photon ring diameters with DM parameters, to be robust across different emission profiles.
Model 1 is designed to represent a standard, geometrically thin, optically thick accretion disk, analogous to the classical Novikov-Thorne model. The emissivity is sharply peaked at the innermost stable circular orbit (ISCO), which is the radius where gravitational energy is most efficiently released, and follows a rapid power-law decay outwards. This profile is characteristic of radiatively efficient accretion scenarios, often seen in luminous active galactic nuclei (AGN).
@@
Model 2 serves as an exploratory case to probe the observational appearance when the emission peak is located extremely close to the unstable photon orbits at $r_{p}$. While not a standard astrophysical scenario, this profile could hypothetically arise from localized energy release, such as magnetic reconnection events occurring near the photon sphere. Its primary purpose is to test the most extreme gravitational lensing effects on an emission source situated at the boundary of photon capture.
@@
Model 3 is chosen to represent a physically distinct scenario: a geometrically thick, optically thin, and radiatively inefficient accretion flow, such as an Advection-Dominated Accretion Flow (ADAF). This type of flow is believed to be present in low-luminosity AGN like $\text{M87}^{*}$ and $\text{SgrA}^{*}$. The emissivity profile is consequently broader, with a more gradual decay, reflecting emission that is spread over a much wider radial range rather than being concentrated at the inner edge.
@@ IV Shadows with spherical accretions
Having analyzed thin disk accretion, we now turn our attention to spherical accretion flows, which provide alternative astrophysical scenarios. We will investigate both static and infalling accretion models to determine how the Hernquist DM halo impacts the corresponding observational signatures.
@@ IV.2 The infalling spherical accretion
Similarly, we maintain the assumption that the emission is monochromatic. Thus, for the infalling accrection, the observed intensity takes the form
@@ V Conclusion
In this work, we have conducted a thorough investigation of the optical appearance and shadows of a Schwarzschild black hole embedded within a Hernquist DM halo. Our findings demonstrate that the presence of a Hernquist DM halo substantially alters the observable characteristics of the black hole shadow across a variety of accretion scenarios. In particular, the photon sphere expands significantly as $\rho_{c}M^{2}$ and $r_{s}$ increase, as seen in Figs.7-12. Furthermore, we find that the core radius $r_{s}$ of the DM halo influences the growth of the photon sphere more significantly than the central density $\rho_{c}$. We use the Schwarzschild black hole with $b_{p}^{\text{Sch}}=3\sqrt{3}M$ as a reference caseLuminet (1979); Gralla et al. (2019). For the parameter set $\rho_{c}M^{2}=0.4$ and $r_{s}=0.2$, the shadow diameter increases relative to the Schwarzschild case by $\Delta\theta/\theta_{\text{Sch}}=\left(b_{p}^{\text{SH}}-b_{p}^{\text{Sch}}\right)/b_{p}^{\text{Sch}}=1.89\%$ Perlick and Tsupko (2022). Keeping $r_{s}=0.2$, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ rises to $3.78\%$ at $\rho_{c}M^{2}=0.8$. Notably, if we increase $r_{s}/M$ from $0.2$ to $0.4$ while keeping $\rho_{c}M^{2}=0.4$ fixed, the relative change climbs markedly to $14.40\%$. In the limiting case, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ reaches $29.15\%$ when $\rho_{c}M^{2}=0.8$ and $r_{s}=0.4$. However, as shown in Figs. 10 and 12, the peak observed intensity is significantly diminished, resulting in a substantially dimmer overall image. The geometric expansion of the photon trajectory is controlled by the Hernquist DM halo’s gravitational impact, although radiative transfer processes like gravitational redshift and Doppler de-boosting are the main causes of the intensity suppression. Our calculations indicate that while the presence of a Hernquist dark matter halo systematically increases the size of the shadow and the emission ring, the dynamical state of the accretion flow (static versus infalling) has a comparable, if not more significant, influence on the absolute level and distribution of brightness. This suggests that, when using EHT observations to constrain the central dark matter distribution, it is imperative to have a prior assumption or an independent constraint on the physical state of the accretion flow.
@@
  • Exclusion Region: The expected shadow deviation approaches $30\%$ in high-density areas. The probability of such high-density Hernquist DM regions in the central region of the Milky Way is essentially ruled out by this large signal, which is much larger than current observational errors of $\sim 10\%$.
@@
Importantly, our research offers significant insights into potential observable features in real astrophysical systems like $\text{M87}^{*}$, which are typically described by the Kerr metric Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023), even though our model is based on a static, modified Schwarzschild background. The predicted variation in shadow size, which manifests as a positive correlation between the photon ring diameter and Hernquist DM parameters, serves as a valuable theoretical framework for probing the DM distribution around supermassive black holes. The current EHT observation for $\text{M87}^{*}$ is approximately $42\pm 3\,\mathrm{\mu as}$, whose relative uncertainty is about $7.14\%$ Collaboration (2019a), which is consistent with the Kerr vacuum solution. Our model provides a theoretical prediction that may be tested by current or next-generation EHT measurements, since this stands in stark contrast to the $\sim 2\%$ to $\sim 30\%$ rise anticipated by Hernquist DM halos. Consequently, when an anomalously large shadow detected in the future indicate the presence of a Hernquist DM halo, our current analysis suggests that the absence of significant shadow enlargement already places stringent constraints on the parameter space of viable Hernquist DM models.

### v4

# Probing Hernquist dark matter through the optical appearance of black holes: A comprehensive study of various accretions

@@ Abstract
The observational appearance of a black hole is critically dependent on the surrounding accreting matter, in particular on the central brightness depression and photon ring structure. We perform a systematic comparative analysis of the observational signatures of a Schwarzschild black hole embedded in a Hernquist dark matter (DM) halo under three distinct accretion scenarios: a geometrically thin disk, a static spherical flow, and an infalling spherical flow. For the thin disk model, we find that direct emission dominates the total observed intensity, while the size and brightness of the lensing and photon rings serve as sensitive probes of the Hernquist DM parameters. From a geometric perspective, the Hernquist DM halo significantly enlarges the photon sphere, resulting in an observable critical curve radius approximately $2\%$ to $30\%$ larger than in the vacuum case. Regarding the radiative signatures, the measured intensity profiles, which rely on the particular accretion models, show a general brightness suppression, which is especially affected by the Doppler de-boosting in the infalling scenario. Our results suggest that the size of the central brightness depression and the brightness profile of the black hole image provide a valuable theoretical framework for constraining the distribution of dark matter in galactic centers.
@@ I Introduction
To investigate the observational characteristics of black hole images and search for potential indications of novel physics, it is essential to employ simplified accretion models. Schwarzschild black holes were studied for the spherically symmetric accretion model, another accretion scenario, and it was shown that under such accretion settings the shadow has robust properties, where its shape and size are determined by spacetime geometry rather than accretion details Gralla et al. (2019); Tsukamoto et al. (2014); Tsukamoto (2018). Conversely, investigations of optically thin and geometrically thin accretion models show that the appearance of the central brightness depression is strongly related to the position of the emission source Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021b); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Direct emission, lensing rings, and photon rings are the three main components that these models usually appear as in the seen image. These distinct signatures offer a potential means to distinguish GR black holes from other compact objects or black holes in modified gravity theories via observational data Zeng and Zhang (2020); Zeng et al. (2020); Li and He (2021b); He et al. (2022); Peng et al. (2021); Zeng et al. (2022, 2025). Consequently, EHT observations of black hole shadows are crucial for constraining parameters in modified gravity theories and testing fundamental physics Shaikh et al. (2019); Davoudiasl and Denton (2019); Roy and Yajnik (2020).
@@
In this paper, we present a systematic comparative study of the shadow of Schwarzschild-Hernquist black holes with different accretion models. In order to clarify the associated impacts of Hernquist DM halo parameters on the photon sphere and observed brightness, we thoroughly investigated three thin disk accretion models and two spherical accretion models. Our results demonstrate that different accretion regimes produce distinctly different images under the same background geometry. Specifically, pictures made up of the central brightness depression, lensing ring, and photon ring are created via thin disk accretion. The shadow and photon ring are spherically symmetric for spherical accretion. A darker shadow than under static conditions results from the extra Doppler effect introduced by falling accretion Li and He (2021a, b); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). We aim to close a gap in the references on Hernquist metrics and accretion models by revealing scaling rules for shadow size and brightness as functions of the Hernquist DM halo distribution. This offers a potential diagnostic framework for constraining Hernquist DM characteristics.
@@
Before proceeding, we clarify the terminology adopted in this work. The critical curve refers to the projection of the photon sphere onto the observer’s image plane, characterized by the critical impact parameter $b_{p}$. The term shadow, in its classical sense following Falcke Falcke et al. (2000), denotes the region enclosed by the critical curve from which no photons reach the distant observer. However, as emphasized by Gralla Gralla et al. (2019), the observed central brightness depression does not necessarily coincide with the critical curve and depends on the properties of the accretion flow Falcke et al. (2000); Gralla et al. (2019); Macedo et al. (2024); Feng and Zhang (2026). For the spherical accretion models, the boundary of the central dark area closely traces the critical curve, so the classical notion of the shadow applies directly. In contrast, for the thin disk models, the observed dark region is shaped by the emission profile and may differ from the classical shadow.
@@ II Light deflection and photon orbit
Here, the Newton constant $G$ has been set to 11 for simplicity. Naturally, this black hole metric (10) reduces to the Schwarzschild black hole in the limit $\rho_{c}\to 0$ or $r_{s}\to 0$. In Fig.1, we depict the behavior of the metric function $f(r)$ to show how the Hernquist DM halo affects spacetime. The largest root of the equation $f(r_{h})=0$ defines the event horizon radius $r_{h}$. This critical point in the Schwarzschild scenario is located at $r_{h}=2M$. However, the DM halo causes the curve to cross the zero axis at a wider radius, as seen in Fig.1. In particular, a monotonic expansion of the event horizon is induced by an increase in either the core radius $r_{s}$ or the central density $\rho_{c}$. This geometric change indicates that the input from the DM halo may successfully boost the black hole’s gravitational capture ability.
[…]
@@
It is important to note that the parameter space explored in this work extends beyond the currently observationally favored region. The EHT observations of M87 constrain the angular shadow size to approximately $42\pm 3\,\mathrm{\mu as}$ Collaboration (2019a), corresponding to a critical impact parameter $b_{p}/M_{\text{BH}}\lesssim 5.56$ Macedo et al. (2024); Feng and Zhang (2026). From Tab. 1, the configurations with $r_{s}/M=0.2$ (yielding $b_{p}/M\approx 5.29$$5.39$) remain compatible with this constraint. However, configurations with $r_{s}/M=0.4$ produce significantly larger values ($b_{p}/M\approx 5.94$$6.71$), exceeding the EHT bound. In particular, the combination $\rho_{c}M^{2}=0.8$ and $r_{s}/M=0.4$ yields a shadow enlargement of $\sim 29\%$ relative to the Schwarzschild case, well above the current observational uncertainty of $\sim 10\%$. Ref. Feng and Zhang (2026) has derived an upper bound on the halo compactness of $\mathcal{C}\leq 0.092$ from the EHT data. We emphasize that the inclusion of these large-parameter configurations is intentional: they serve as a phenomenological exploration of the theoretical upper bounds on the Hernquist DM effects, rather than as claims of astrophysical viability. A detailed assessment of the observational constraints is provided in Sec. V.
@@
Figure 3: The behaviour of light trajectories under various $\rho_{c}$ and $r_{s}$ in the polar coordinate system $(r,\varphi)$. The solid black disk and the green dashed circle indicates the event horizon and the photon sphere, respectively. The yellow lines correspond to scattered rays $(b>b_{p})$, the blue lines to captured light $(b<b_{p})$ and the red one means the critical unstable orbit $(b=b_{p})$.
@@ III Observational features of thin disk emission
In astrophysical settings, black holes are typically surrounded by large amounts of accreting matter. This makes the study of the observational appearance of Schwarzschild-Hernquist black holes particularly relevant from an observational perspective. We note that in the thin disk scenario discussed below, the boundary of the observed central brightness depression is determined by the emission profile and does not necessarily coincide with the critical curve at $b=b_{p}$, in contrast to the spherical accretion case. This distinction has been highlighted in recent works Gralla et al. (2019); Gralla and Lupsasca (2020); Macedo et al. (2024); Paugnat et al. (2022).In this section we will investigate the central brightness depression and ring structure of a Schwarzschild-Hernquist black hole surrounded by a disk-like accretion disk that is optically and geometrically thin.
@@ III.1 Direct emission, lensing ring and photon ring
Figure 5: Photon behaviour as a function of impact parameter $b$ in the Schwarzschild-Hernquist black hole. The blue lines represent trajectories contributing to direct emission ($n<3/4$), yellow lines constitute the lensing ring ($3/4<n<5/4$), and red lines mean the photon ring ($n>5/4$). The black disk and the green dashed circle indicates the event horizon and the photon sphere $r_{p}$, respectively.
@@
In favour of the number of orbits $n(b)$, the emission can be categorized into three distinct zones Gralla et al. (2019); Gralla and Lupsasca (2020); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025); Li and He (2021a, b). These are lensing rings, photon rings, and direct emission rings corresponding to $n<3/4$, $3/4<n<5/4$, and $n>5/4$, respectively. These zones correspond to light rays that intersect the equatorial plane once, twice, or more than twice, respectively Gralla et al. (2019); Gralla and Lupsasca (2020). For various values of the parameters $\rho_{c}$ and $r_{s}$, Fig.4 shows the functional dependence of $n(b)$ on the impact parameter $b$. The domains of direct emissions, lensing rings, and photon rings are shown by different colors. We have indicated the ring and the direct emission locations in Fig.5. The function $n(b)$ diverges as bapproaches the critical value $b_{p}$, indicating the existence of unstable photon orbits, as shown in Fig.4. The curves for larger DM parameters $\rho_{c}$ and $r_{s}$ are shifted to the right, consistent with the increase in $b_{p}$ reported in Tab.1. Tab.2 lists the ranges of the impact parameter bcorresponding to direct emission, the lensing ring, and the photon ring for various values of $\rho_{c}$ and $r_{s}$. It is evident from the data that the range of the impact parameter $b$ for all emission types increases as the parameters increase. Furthermore, the total number of photon orbits $n(b)$ will peak inside a small area when the impact parameter $b$ gets closer to the critical condition $b\sim b_{p}$. The extremely narrow ranges for the photon ring listed in Tab.2 correspond to the thin red band visible in Fig.5. The angular radius of the critical curve (determined by $b_{p}$) also increases with larger parameters. However, this enlargement does not directly translate to an increase in the brightness of the ring; the observed intensity depends on the complex interplay between the emission profile and the light bending.
We remark that an alternative classification scheme, widely adopted in recent studies Gralla and Lupsasca (2020); Johnson et al. (2020); Cárdenas-Avendaño and Lupsasca (2023); Paugnat et al. (2022); Vincent et al. (2022); Macedo et al. (2024), labels photon trajectories by the number of half-turns $m$ performed around the black hole, rather than the number of orbits $n$ used in the present work. In this notation, $m=0$ corresponds to direct emission, while $m=1,2,\ldots$ denote successive photon rings. The correspondence with our classification is straightforward: direct emission ($n<3/4$) maps to $m=0$, the lensing ring ($3/4<n<5/4$) to $m=1$, and the photon ring ($n>5/4$) to $m=2$. We retain the notation of Ref. Gralla et al. (2019) throughout this paper for internal consistency.
@@ III.3 Observational features
To compute the observational appearance of the thin disk, we employ three phenomenological emissivity models Wang et al. (2022); Yang et al. (2023); Li and He (2021b); Zeng and Zhang (2020); Zeng et al. (2020, 2022, 2025). These three emission profiles follow the toy-model introduced by Gralla Gralla et al. (2019), which has been widely adopted in the black hole shadow literature. It is important to note that these models are simplified representations and are strategically chosen to represent a diverse range of plausible astrophysical scenarios from efficient, geometrically thin disks to inefficient, thick flows Meng et al. (2023); Gan et al. (2021); Guo et al. (2022); Chen et al. (2022). This approach allows us to systematically explore how the observational signatures of the Hernquist DM halo depend on the spatial distribution of the emitting matter. While the precise quantitative features of the image, such as peak brightness, are model-dependent, we expect the qualitative findings, particularly the scaling of the critical curve and photon ring diameters with DM parameters, to be robust across different emission profiles.
Model 1 is designed to represent a standard, geometrically thin, optically thick accretion disk, inspired by the classical Novikov-Thorne model Shakura and Sunyaev (1973); Novikov and Thorne (1973); Page and Thorne (1974). The emissivity is sharply peaked at the innermost stable circular orbit (ISCO), which is the radius where gravitational energy is most efficiently released, and follows a rapid power-law decay outwards. This profile is characteristic of radiatively efficient accretion scenarios, often seen in luminous active galactic nuclei (AGN) Yuan and Narayan (2014); Koratkar and Blaes (1999).
@@
Model 2 serves as an exploratory case to probe the observational appearance when the emission peak is located extremely close to the unstable photon orbits at $r_{p}$. While not a standard astrophysical scenario, this profile could hypothetically arise from localized energy release, such as magnetic reconnection events occurring in the innermost black hole magnetosphere Ripperda et al. (2022); Sironi and Beloborodov (2020). Its primary purpose is to test the most extreme gravitational lensing effects on an emission source situated at the boundary of photon capture.
@@
Model 3 is chosen to represent a physically distinct scenario: a geometrically thick, optically thin, and radiatively inefficient accretion flow, such as an Advection-Dominated Accretion Flow (ADAF) Narayan and Yi (1994); Narayan (1995); Yuan and Narayan (2014). This type of flow is believed to be present in low-luminosity AGN like $\text{M87}^{*}$ and $\text{SgrA}^{*}$ Yuan et al. (2003); Collaboration (2019a). The emissivity profile is consequently broader, with a more gradual decay, reflecting emission that is spread over a much wider radial range rather than being concentrated at the inner edge.
@@ IV Shadows with spherical accretions
Having analyzed thin disk accretion, we now turn our attention to spherical accretion flows, which provide alternative astrophysical scenarios. We will investigate both static and infalling accretion models to determine how the Hernquist DM halo impacts the corresponding observational signatures. Unlike the thin disk case, for spherically symmetric accretion the shadow boundary is determined by the critical curve, and the classical notion of the black hole shadow applies directly Narayan et al. (2019); Falcke et al. (2000).
@@ IV.2 The infalling spherical accretion
Similarly, we maintain the assumption that the emission is monochromatic. Thus, for the infalling accretion, the observed intensity takes the form
@@ V Conclusion
In this work, we have conducted a thorough investigation of the optical appearance and shadow properties of a Schwarzschild black hole embedded within a Hernquist DM halo. Our findings demonstrate that the presence of a Hernquist DM halo substantially alters the observable characteristics of the black hole image across a variety of accretion scenarios. In particular, the photon sphere expands significantly as $\rho_{c}M^{2}$ and $r_{s}$ increase, as seen in Figs.7-12. Furthermore, we find that the core radius $r_{s}$ of the DM halo influences the growth of the photon sphere more significantly than the central density $\rho_{c}$. We use the Schwarzschild black hole with $b_{p}^{\text{Sch}}=3\sqrt{3}M$ as a reference caseLuminet (1979); Gralla et al. (2019). For the parameter set $\rho_{c}M^{2}=0.4$ and $r_{s}=0.2$, the critical impact parameter $b_{p}$ increases relative to the Schwarzschild case by $\Delta\theta/\theta_{\text{Sch}}=\left(b_{p}^{\text{SH}}-b_{p}^{\text{Sch}}\right)/b_{p}^{\text{Sch}}=1.89\%$ Perlick and Tsupko (2022). Keeping $r_{s}=0.2$, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ rises to $3.78\%$ at $\rho_{c}M^{2}=0.8$. Notably, if we increase $r_{s}/M$ from $0.2$ to $0.4$ while keeping $\rho_{c}M^{2}=0.4$ fixed, the relative change climbs markedly to $14.40\%$. In the limiting case, the relative change rate $\Delta\theta/\theta_{\text{Sch}}$ reaches $29.15\%$ when $\rho_{c}M^{2}=0.8$ and $r_{s}=0.4$. However, as shown in Figs. 10 and 12, the peak observed intensity is significantly diminished, resulting in a substantially dimmer overall image. The geometric expansion of the photon trajectory is controlled by the Hernquist DM halo’s gravitational impact, although radiative transfer processes like gravitational redshift and Doppler de-boosting are the main causes of the intensity suppression. Our calculations indicate that while the presence of a Hernquist dark matter halo systematically increases the critical impact parameter and the emission ring radius, the dynamical state of the accretion flow (static versus infalling) has a comparable, if not more significant, influence on the absolute level and distribution of brightness. This suggests that, when using EHT observations to constrain the central dark matter distribution, it is imperative to have a prior assumption or an independent constraint on the physical state of the accretion flow.
@@
  • Exclusion Region: The expected deviation of $b_{p}$ approaches $30\%$ in high-density areas. The probability of such high-density Hernquist DM regions in the central region of the Milky Way is essentially ruled out by this large signal, which is much larger than current observational errors of $\sim 10\%$.
@@
Importantly, our research offers significant insights into potential observable features in real astrophysical systems like $\text{M87}^{*}$, which are typically described by the Kerr metric Collaboration (2019a, b, c, d, e, f, 2021a, 2021b, 2023), even though our model is based on a static, modified Schwarzschild background. The predicted variation in the critical impact parameter size, which manifests as a positive correlation between the photon ring diameter and Hernquist DM parameters, serves as a valuable theoretical framework for probing the DM distribution around supermassive black holes. The current EHT observation for $\text{M87}^{*}$ is approximately $42\pm 3\,\mathrm{\mu as}$, whose relative uncertainty is about $7.14\%$ Collaboration (2019a), which is consistent with the Kerr vacuum solution. Our model provides a theoretical prediction that may be tested by current or next-generation EHT measurements, since this stands in stark contrast to the $\sim 2\%$ to $\sim 30\%$ rise anticipated by Hernquist DM halos. Consequently, when an anomalously large shadow detected in the future indicate the presence of a Hernquist DM halo, our current analysis suggests that the absence of significant shadow enlargement already places stringent constraints on the parameter space of viable Hernquist DM models.
It is instructive to compare our findings with related studies of dark matter effects on black hole imaging. Ref. Macedo et al. (2024) investigated the optical appearance of black holes surrounded by a Hernquist-type dark matter halo using the exact non-perturbative solution reported in Ref. Cardoso et al. (2022). Their analysis combined hot-spot astrometry and the Gralla-Lupsasca-Marrone emission models. For configurations saturating the EHT shadow-size constraint with compactness $\mathcal{C}\sim 1/15$, they found that the photon ring locations shift outward by approximately $7\%$ and their widths increase by up to $\sim 25\%$ compared to Schwarzschild case, while the relative luminosities of successive photon rings are only marginally affected. Their overall conclusion is that the optical appearance of the dark matter halo black hole closely mimics that of a Schwarzschild black hole under current observational precision. This is consistent with our finding that, for the lower-parameter configurations with $\rho_{c}M^{2}=0.4$ and $r_{s}/M=0.2$, the critical curve enlargement is merely $\sim 2\%$, falling below the EHT detection threshold. More recently, Ref. Feng and Zhang (2026) conducted a comprehensive study of the shadow and quasi-normal modes of the Schwarzschild-Hernquist black hole across a broad parameter space. They derived an upper bound on the halo compactness $\mathcal{C}\leq 0.092$ from EHT data and obtained an analytical formula for the QNM frequency redshift which is valid for $\mathcal{C}\leq 0.3$. This QNM redshift is driven by the same gravitational potential deepening that causes the shadow enlargement and brightness suppression reported in our work, and provides a complementary gravitational-wave channel for constraining the Hernquist DM halo parameters. Our work extends these previous studies by providing a systematic comparison across three distinct accretion scenarios, namely thin disk, static spherical, and infalling spherical accretion. This comparison reveals that while the geometric signature of the critical curve enlargement is robust across all models, the dynamical state of the accretion flow has a comparable, if not greater, influence on the brightness distribution. This model-dependent radiative behavior underscores the importance of incorporating realistic accretion physics when using black hole images to constrain dark matter distributions.